OpenStudy (anonymous):

given f(x)=(x^2)-2x a: Find f ' (x) b: Find the slopes of the lines tangent to the graph of f at x=-2, -1, and 2 c: Graph f, and sketch in the tangent lines at x=-2, -1, and 2

OpenStudy (anonymous):

a) Use power rule. b) It asks you to find f'(-2), f'(-1), and f'(2).

OpenStudy (anonymous):

I'm taking and online course and am having trouble understanding how to do those things, thats my problem =/

OpenStudy (anonymous):

2x-2? for A?

jimthompson5910 (jim_thompson5910):

a) like RD⁴² is saying you use the power rule so if f(x) = x^3 for example, then f ' (x) = 3x^2 in general, if f(x) = x^n, then f ' (x) = n*x^(n - 1)

jimthompson5910 (jim_thompson5910):

so good, you would derive f(x) = x^2 - 2x to get f ' (x) = 2x - 2

jimthompson5910 (jim_thompson5910):

just saw your response lol

jimthompson5910 (jim_thompson5910):

now you would use this for part b

jimthompson5910 (jim_thompson5910):

f ' (x) = 2x - 2 f ' (-2) = 2(-2) - 2 f ' (-2) = -4 - 2 f ' (-2) = -6 This means that the slope of the tangent line at x = -2 (on f(x)) is -6

jimthompson5910 (jim_thompson5910):

you would repeat for x = -1 and x = 2

OpenStudy (anonymous):

Okay! Formulas are one thing but putting them on a graph?! Theres a real issue haha! You are an amazing person thanks so much for taking the time! I wish you were a tutor!

jimthompson5910 (jim_thompson5910):

lol I am a tutor (sorta) to graph f(x)=(x^2)-2x, you can plot a bunch of points or you can use a graphing calculator

OpenStudy (anonymous):

Well an online tutor because I'd seriously pay you, I actually understand things when you go through them with me!

jimthompson5910 (jim_thompson5910):

here is the graph of f(x)=x^2-2x shown in red

jimthompson5910 (jim_thompson5910):

now locate x = -2 and plot the point (-2, 8) on the graph

jimthompson5910 (jim_thompson5910):

you can see that when x = -2, y is... y = x^2 - 2x y = (-2)^2 - 2(-2) y = 4 + 4 y = 8 So that algebraically explains why the point (-2, 8) lies on the graph

jimthompson5910 (jim_thompson5910):

from there, you draw a tangent line that goes through (-2, 8) you do this by drawing a line with a slope of -6 that goes through (-2, 8) so start at (-2, 8) and go down 6 and to the right 1 to get to the next point (-1, 2) the added green line is the tangent line at x = -2

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

It really does...like I said, you're awesome! And you dont make me feel like an idiot!

jimthompson5910 (jim_thompson5910):

and I do offer paid tutoring, but tutoring I do on here is free since this site is free

jimthompson5910 (jim_thompson5910):

sweet, glad it's making more sense and no definitely not, you're learning calculus here lol

OpenStudy (anonymous):

Wheat site do you offer paid tutoring and how much?

jimthompson5910 (jim_thompson5910):

I just offer it through email or chat no official site really (at the moment I guess)

OpenStudy (anonymous):

If you have any availability I am really interested!

jimthompson5910 (jim_thompson5910):

let me know if you still need help on this problem, but for the most part it sounds like you got this one (which is great)